The frequency of light in a material is 4*10^14Hertz and wave length is 5000amstrong. The refractive index of the material will be?

Grade:10

1 Answers

sathya
6 years ago
 
frequency=velocity/wave lenght
so 4*10^14=X/5000*10^-10(where X=VELOCITY)
X=2*10^8m/s
refractive index =light speed/velocity
Y=3*10^8/2*10^8(where Yis refractive index)
Y=1.5